OpenStudy (anonymous):

Please help me :( For the functions f(x) = 2x + 3 and g(x) = 6x + 2, which composition produces the greatest output? (5 points) Neither composition produces an output. Both compositions produce the same output. f(g(x)) produces the greatest output. g(f(x)) produces the greatest output.

OpenStudy (anonymous):

f(x)= 2x+3 g(x)=6x+2 f(g(x))= f(6x+2)= 2(6x+2)+3= 12x+7 g(f(x))= g(2x+3)=6(2x+3)+2= 12x+20 Regardless of the value of x, adding 20 to the 12x will produce a larger outcome than adding 7 to 12x. g(f(x)) produces the greatest output